OpenStudy (anonymous):

3x + 2y = 4 2x - y = 5 This system of equations has no solution. has one solution. is coincident.

OpenStudy (anonymous):

one second @xo_kansasprincess_xo

OpenStudy (anonymous):

Since 2x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2x from both sides. 3x+2y=4 −y=−2x+5 Divide each term in the equation by −1. 3x+2y=4 y=2x−5 Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is 2x−5. 3x+2(2x−5)=4 y=2x−5 Simplify each term. 3x+4x−10=4 y=2x−5 Add 3x and 4x to get 7x. 7x−10=4 y=2x−5 Since −10 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 10 to both sides. 7x=10+4 y=2x−5 Add 10 and 4 to get 14. 7x=14 y=2x−5 Divide each term in the equation by 7. x=2 y=2x−5 Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is 2. x=2 y=2(2)−5 Multiply 2 by 2 to get 4. x=2 y=4−5 Subtract 5 from 4 to get −1. x=2 y=−1 The solution to the system of equations can be represented as a point. (2,−1)

OpenStudy (anonymous):

therefore it has on solution

OpenStudy (anonymous):

8x + 13y = 2 y = x - 2 Solve the system of equations

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

(4/3, -2/3) (4/21, -38/21) No solution

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 3x + 2y = 4 }\) \(\large\color{black}{ \displaystyle 2x - y = 5 }\) \(\large\color{black}{ \displaystyle 2x - y = 5~~~~\Rightarrow~~~~y=2x-5}\) \(\large\color{black}{ \displaystyle 3x + 2y = 4~~~~\Rightarrow~~~~3x + 2(2x-5) = 4 }\)

OpenStudy (solomonzelman):

don't know what question marks say, I guess I am out.

OpenStudy (anonymous):

What?

OpenStudy (solomonzelman):

if you don't understand my post, then don't worry about it....

OpenStudy (anonymous):

Did I say I didn't understand it? Lol no I didn't. So don't get an attitude with me. ;)